OpenStudy (anonymous):

Express 12.6 as a fraction

Directrix (directrix):

12.6 is 12 and 6/10. There are ten tenths in one. In twelve, there are 120 tenths. 120/10 + 6/10 = 126/10 = 63/5

OpenStudy (anonymous):